OpenStudy (anonymous):

Algebra 2 help! D:

OpenStudy (anonymous):

Is this correct?

OpenStudy (anonymous):

i got x=-4,3 not sure if thats right however

OpenStudy (anonymous):

-4.3?

OpenStudy (anonymous):

-4 and 3

OpenStudy (anonymous):

Ohhh.. well it's only one answer..

OpenStudy (phi):

First, you can check your answer 5.27 by using that value for x in the equation you will see you get different numbers on the each side of the equation. that means 5.27 is not the answer

OpenStudy (phi):

to solve, square both sides \[ (x+2)^2 = (\sqrt{3x+16})^2 \\ x^2 +4x+4= 3x+16 \]

OpenStudy (anonymous):

I did that part.

OpenStudy (phi):

you should simplify to get \[ x^2 +x -12=0 \]

OpenStudy (anonymous):

Yes

OpenStudy (phi):

that factors

OpenStudy (anonymous):

Yes with the quadratic function.

OpenStudy (phi):

you mean quadratic formula. Though the quadratic formula works, you don't need it. factor into (x+4)(x-3)= 0 that means either x+4=0 or x-3 =0 x=-4 and x= 3 are solutions (as already pointed out by lewopz)

OpenStudy (phi):

notice that if you test x=3 it works. also notice that if you test x=-4, it will not work (you get the -2=2 which is not true) the x=-4 is an extraneous solution

OpenStudy (anonymous):

So I would put 3 as the answer..

OpenStudy (phi):

if you use the quadratic formula you should get \[ \frac{-1±\sqrt{1+4\cdot 12}}{2} = \frac{-1±\sqrt{1+48}}{2} \\ = \frac{-1±\sqrt{49}}{2}\\ = \frac{-1±7}{2} = -4 \text{ or } 3 \]

OpenStudy (phi):

try out your answer x=3 on the left side you get x+2 becomes 3+2 = 5 on the right side \[ \sqrt{3x+16}= \sqrt{3\cdot3 +16} = \sqrt{9+16}= \sqrt{25}=5\] both sides are the same. so 3 works.